OpenStudy (anonymous):

Marlene is a sales associate at a department store. Each week she earns $100 plus a 5% commission of all her sales. This week she earned $160. How much, in dollars, were Marlene's sales this week? (Do not put a $ in your answer.) Please help @Mertsj

OpenStudy (anonymous):

@Ishaan94 Help?

OpenStudy (anonymous):

Anyone? :(

OpenStudy (anonymous):

@jim_thompson5910 ???

OpenStudy (anonymous):

I she makes $100 plus 5% commission and she made $160 then you take the $60 times the 5% and the answer would be 3

OpenStudy (anonymous):

I don't think that's quite correct thank thank you though. :)

OpenStudy (anonymous):

168

OpenStudy (anonymous):

what is 60 x 5...

OpenStudy (anonymous):

its definetly not 168.

OpenStudy (anonymous):

it wouldn't be 5 it would be .05

OpenStudy (anonymous):

your talking % not a whole 5

OpenStudy (anonymous):

But we are trying to find out how much she sold..not the percentage of how much. we already know the percentage.

OpenStudy (anonymous):

but you have to take the total multiply but 5% or .05 then add it to your whole number

OpenStudy (anonymous):

that she got herself.

jimthompson5910 (jim_thompson5910):

100+0.05x = 160 0.05x = 160-100 0.05x = 60 x = 60/0.05 x = 1200 So she made $1200 in sales

OpenStudy (anonymous):

THANK YOU!!!! <3

jimthompson5910 (jim_thompson5910):

yw